OpenStudy (anonymous):

When solving a system of equations by the process of elimination, if both variables get canceled as soon as I alter one of the equations, how do I figure out the value of each variable? This is the problem I'm stuck on: 2x + 4y = -1 4x + 8y = 0 I tried multiplying the first equation by -2, which results in both variables being canceled out. PLEASE HELP!

OpenStudy (anonymous):

I think it means there are no solutions.

OpenStudy (anonymous):

Ok, yes. Write them both in \(y = mx + b\) form. \[y = -\frac{1}{2}x - \frac{1}{4}\]\[y = -\frac{1}{2}x\] Those lines are parallel. No solutions.

OpenStudy (anonymous):

Thank you so much! I was losing it over this one problem, lol.